OpenStudy (anonymous):

For example, there are balls marked b1, b2, up to b15; i16, i17, up to i30; and so on. Assuming one bingo ball is selected at random, determine the odd against it containing the letter G. Total G balls =15 Total Non-G balls=60 Total number of balls=75 60/75 15/75 60/75*15/75=60/15 60/15=4/1 Odds against it being the G ball =4:1

OpenStudy (anonymous):

I"m trying to find the probability of NOT containing the letter G

OpenStudy (anonymous):

This is what I have so Far Total G balls =15 Total Non-G balls=60 Total number of balls=75 60/75 15/75 60/75*15/75=60/15 60/15=4/1 Odds against it being the G ball =4:1

OpenStudy (anonymous):

Any tips would be very helpful. Thank you

OpenStudy (anonymous):

# of G-balls: 15 ----- # of non-G balls: 60

OpenStudy (anonymous):

P(non-G) = 60/75 P(G-balls) = 15/75 --- Odds against it being a G-ball = (60/75)/(15/75) = 60/15 = 4:1